The next four questions refer to the following choices: a. Rate k [M] b. Rate k [M][N] c. Rate k[M][N] d. Rate k[M] [N] e. Rate k[N] 39. Doubling the concentration of M has no effect on the reaction rate 40. Doubling the concentration of M and Nincreases reaction rate by 2 41. Doubling the concentration of M only quadruples the reaction rate 42. Doubling the concentration of M and N increases the reaction rate by eightfold. Please answer 39-42!!!
Answer #1

Q39.

if we double M, there should no be effect, since it is not present in th eRATE equation, this must be "E" since it is the only which does not contain M

Q40

if we doulbe M and N, then rate increases by 2, therefore, only A is true, since doubling M will increase rate by 2

Q41

if we doulb eM only it will increase by 2, then M must be to the second power

so

the best answer is D

Q42

If we doulbe M and N, then there is a 8x increase

Best answer msut be n^2 n^1 = 2^2 * 2^1 = 4*2 = 8

that is, C